505g of solid lead at 100 C is heated to its melting point and converted completely to liquid lead. Calculate the amount of thermal energy required for this to happen.
1
Expert's answer
2020-02-26T10:09:14-0500

"m=505 \ g= 0.505 \ kg"

"T_i=100^{\circ} C"

"T_f=327.5^{\circ}C"

"c=130 \ J(^{\circ}C \ kg)"

"L_{f}=23\times 10^3 \ J/kg"


"Q \ - ?"


Solution:

At first, we need to increase the temperature of the lead from 100ºC to 327.5ºC (melting point).

"Q_1=cm(T_{f}-T_{i})=cm\Delta T"

"Q_1=130\times 0.505\times 227.5=14935.375 \ J"

Heat necessary for melting the lead:

"Q_2=L_{f} m"

"Q_2=23\times 10^3\times 0.505=11615 \ J"

Total heat required for this to happen:

"Q=Q_1+Q_2=26550.375\ J=26.5 \ kJ"


Answer: "26.5 \ kJ."